# Material Point Methods on Unstructured Tessellations: A Stable Kernel Approach With Continuous Gradient Reconstruction

### v1

@@ Abstract
The Material Point Method (MPM) is a hybrid Eulerian-Lagrangian simulation technique for solid mechanics with significant deformation. Structured background grids are commonly employed in the standard MPM, but they may give rise to several accuracy problems in handling complex geometries. When using (2D) unstructured triangular or (3D) tetrahedral background elements, however, significant challenges arise (e.g., cell-crossing error). Substantial numerical errors develop due to the inherent $\mathcal{C}^{0}$ continuity property of the interpolation function, which causes discontinuous gradients across element boundaries. Prior efforts in constructing $\mathcal{C}^{1}$ continuous interpolation functions have either not been adapted for unstructured grids or have only been applied to 2D triangular meshes. In this study, an Unstructured Moving Least Squares MPM (UMLS-MPM) is introduced to accommodate 2D and 3D simplex tessellation. The central idea is to incorporate a diminishing function into the sample weights of the MLS kernel, ensuring an analytically continuous velocity gradient estimation. Numerical analyses confirm the method’s capability in mitigating cell crossing inaccuracies and realizing expected convergence.
keywords:
Material Point Method, Unstructured Grid, Cell Crossing Instability, Moving Least Squares
journal: Computer Methods in Applied Mechanics and Engineeringjournal:
@@ 2.1 Governing Equations
In the standard implementation of the MPM, physical quantities are retained at material points and then projected onto background grids for further computation. Equation (6) is discretized within the grid space by leveraging the Finite Element Method (FEM) and then solved using either implicit or explicit time integration schemes. This article focuses on the explicit symplectic Euler time integration method. While the extension to implicit methods is possible and indeed straightforward, that would be orthogonal to the contribution of the article.
@@ 2.2 Explicit MPM Pipeline
The explicit MPM pipeline in each time step is delineated into four main stages: (1) the transfer of material point quantities to the background grid, commonly known as Particle-To-Grid (P2G), (2) the computation of the system’s evolution on this background grid, (3) the back-transfer of the evolved grid quantities to the material points, designated as Grid-To-Particle (G2P), and (4) the execution of supplementary processing tasks, specifically on strain and/or stress to incorporate effects such as elastoplasticity return mapping and material hardening. Finally, the hypothetical deformation incurred on the background grid is reset for the next computational cycle. Algorithm 1 presents an overview of the MPM pipeline, and the main stages are elaborated below.
@@
1:Determine material point-grid connectivity, calculate kernel functions $w_{ip}$
@@
3:          Nodal mass: $m_{i}=\sum_{p}\rho_{p}V_{p}w_{ip}$
4:          Nodal momentum: ${\bm{p}}_{i}=\sum_{p}{\bm{v}}_{p}\rho_{p}V_{p}w_{ip}$
@@
6:          Internal force: ${\bm{f}}_{i}^{\text{int}}=-\sum_{p}V_{p}\bm{\sigma}_{p}\nabla w_{ip}$
7:          Gravity: ${\bm{f}}_{i}^{\text{ext}}=\sum_{p}w_{ip}m_{p}\mathbf{g}_{p}$
@@
14:          Update point velocities: ${\bm{v}}_{p}^{\Delta t}={\bm{v}}_{p}+\Delta t\sum_{i}w_{ip}\ddot{{\bm{x}}}_{i}$
15:          Update point positions: ${\bm{x}}_{p}^{\Delta t}=\sum_{i}w_{ip}\tilde{{\bm{x}}}_{i}$
16:Update deformation gradient: ${\bm{F}}_{p}^{\Delta t}=\left({\bm{I}}+\sum_{i}(\tilde{{\bm{x}}}_{i}-{\bm{x}}_{i})(\nabla w_{ip})^{T}\right){\bm{F}}_{p}$
@@
19:Enforce plasticity, reset grid deformation, advance to next timestep
@@
In the MPM, discrete material points represent physical attributes such as mass, position, and velocity. For a given particle, labeled $p$, and a grid node, labeled $i$, the interpolation function’s value associated with node $i$, evaluated at the spatial position ${\bm{x}}_{p}$ of particle $p$, is represented as $w_{ip}$. Similarly, the gradient of this interpolation function, evaluated at the same location, is denoted as $\nabla w_{ip}$. In the explicit MPM framework, the lumped grid mass is defined as $m_{i}=\sum_{p}\rho_{p}V_{p}w_{ip}$, where $\rho_{p}$ represents the density and $V_{p}$ the volume of each material point. This definition facilitates the momentum calculation on the grid, expressed as
@@
where $\ddot{{\bm{x}}}_{i}$ is the acceleration of grid node $i$. The internal forces ${\bm{f}}^{\text{int}}_{i}$ and external forces ${\bm{f}}^{\text{ext}}_{i}$ acting on the grid nodes are as follows:
${\bm{f}}^{\text{int}}_{i}$ $=-\sum_{p}V_{p}\bm{\sigma}_{p}\nabla w_{ip},$ (8)
${\bm{f}}^{\text{ext}}_{i}$ $=\sum_{p}m_{p}w_{ip}{\bm{b}}_{p}+\sum_{p}m_{p}w_{ip}{\bm{g}}_{p}.$ (9)
The stress tensor $\bm{\sigma}$ is linked to the deformation gradient ${\bm{F}}$ through the constitutive relation
$\bm{\sigma}=C({\bm{F}}),$ (10)
which defines how material deformation influences internal force.
Stage 2: Evolution on the Background Grids
Nodal accelerations are computed using (7). To update the velocities and positions of the grid nodes, a symplectic Euler time integrator is employed:
@@
These equations ensure a time-stepped progression of the grid nodes where the time step size $\Delta t$ is chosen based on the CFL condition.
@@
The FLIP scheme [2] is utilized for all the experiments discussed in Section 3. In FLIP, the material point positions and velocities are updated as
$x_{p}^{\Delta t}$ $=\sum_{i}w_{ip}\tilde{{\bm{x}}}_{i},$ (13)
$v_{p}^{\Delta t}$ $=v_{p}+\Delta t\sum_{i}w_{ip}\ddot{{\bm{x}}}_{i}.$ (14)
Subsequently, the evolution of the deformation gradient ${\bm{F}}$ in (3) is discretized as
${\bm{F}}_{p}^{\Delta t}=\left({\bm{I}}+\sum_{i}(\tilde{{\bm{x}}}_{i}-{\bm{x}}_{i})(\nabla w_{ip})^{T}\right){\bm{F}}_{p}.$ (15)
With an initial ${\bm{F}}^{0}={\bm{I}}$, material point volumes are updated as
@@
Stage 4: Post-Processing and Resetting the Background Grid
This stage encompasses all post-processing tasks such as plasticity return mapping and hardening [46]. In the updated Lagrangian MPM, the grid is reset to a non-deformed state. This can be done by not updating grid positions while discarding other grid information such as velocity and acceleration.
@@ 2.3 Transfer Kernel
In the MPM, the transfer kernel is vital for relaying particle information to adjacent grid nodes. Techniques such as the B-spline MPM [39] and GIMP [36] use a specific compact support function to smoothly influence nearby grid nodes, whereas methods like Moving Least Squares MPM (MLS-MPM) [41] determine the kernel implicitly, based on the proximity of nodes. However, both strategies follow a similar workflow: (1) identifying the set of nearby nodes, and (2) calculating the weights (transfer kernel) for each node in relation to a particle in its influence zone.
This section first introduces the general MLS reconstruction process and the application of MLS-MPM with a comprehensive linear polynomial basis. This is followed by a discussion on a superficially straightforward extension of MLS-MPM to unstructured meshes, highlighting the aforementioned steps of identifying nearby nodes and computing the transfer weights. We then delve into the desirable properties of the kernel, emphasizing why the naive extension fails to yield continuous gradient reconstructions when particles cross cell boundaries. Finally, we offer a solution addressing the issue of discontinuous gradient reconstructions and propose UMLS-MPM.
@@ 2.3.1 Introduction to General MLS and MLS-MPM
The essential concept of Moving Least Squares (MLS) is to approximate a function $u$ at a point ${\bm{z}}$ within the continuous domain surrounding ${\bm{x}}$. This approximation is achieved by employing a polynomial least-squares fit of $u$ based on its sampled values at specific points ${\bm{x}}_{i}$, where each $u_{i}$ denotes the value of $u$ at ${\bm{x}}_{i}$. The functional reconstruction is given by
$u({\bm{z}})={\bm{P}}^{T}({\bm{z}}-{\bm{x}}){\bm{w}}({\bm{x}}),$ (17)
where ${\bm{P}}({\bm{z}}-{\bm{x}})=[p_{0}({\bm{z}}-{\bm{x}}),p_{1}({\bm{z}}-{\bm{x}}),\ldots,p_{l}({\bm{z}}-{\bm{x}})]^{T}$ represents the polynomial basis, while ${\bm{w}}({\bm{x}})=[w_{0}({\bm{x}}),w_{1}({\bm{x}}),\ldots,w_{l}({\bm{x}})]$ comprises the corresponding coefficients, and $l$ indicates the number of polynomial basis functions. The coefficients ${\bm{w}}({\bm{x}})$ are determined by minimizing the cumulative weighted reconstruction errors at the sampled points. This is done by substituting ${\bm{z}}\leftarrow{\bm{x}}_{i}$ into (17) for each sample:
${\bm{w}}({\bm{x}})=\mathop{\rm argmin}\sum_{i\in{\mathcal{B}}_{{\bm{x}}}}d({\bm{x}}_{i}-{\bm{x}})\,||u_{i}-{\bm{P}}^{T}({\bm{x}}_{i}-{\bm{x}}){\bm{w}}({\bm{x}})||^{2}.$ (18)
Here $d$ represents the inverse of a separation function, which is typically a positive value that decreases with increasing separation distance. It acts as the weight for the reconstruction error at each sample. The set ${\mathcal{B}}_{{\bm{x}}}$ encompasses the local region around ${\bm{x}}$ where this weighting function is non-zero; i.e., $d({\bm{x}}_{i}-{\bm{x}})>0$.
This minimization leads to the following solution for ${\bm{w}}$:
${\bm{w}}({\bm{x}})={\bm{M}}^{-1}({\bm{x}}){\bm{b}}({\bm{x}}),$ (19)
where ${\bm{M}}({\bm{x}})=\sum_{i\in{\mathcal{B}}_{{\bm{x}}}}d({\bm{x}}_{i}-{\bm{x}}){\bm{P}}({\bm{x}}_{i}-{\bm{x}}){\bm{P}}^{T}({\bm{x}}_{i}-{\bm{x}})$ and ${\bm{b}}=\sum_{i\in{\mathcal{B}}_{{\bm{x}}}}d({\bm{x}}_{i}-{\bm{x}}){\bm{P}}({\bm{x}}_{i}-{\bm{x}})u_{i}$. Substituting (19) into (17), we obtain the reconstruction
$u({\bm{x}})=\sum_{i\in{\mathcal{B}}_{{\bm{x}}}}d({\bm{x}}_{i}-{\bm{x}}){\bm{P}}^{T}({\bm{z}}-{\bm{x}}){\bm{M}}^{-1}({\bm{x}}){\bm{P}}({\bm{x}}_{i}-{\bm{x}})u_{i}.$ (20)
@@
A special case involves using a complete linear polynomial basis, as is done in MLS-MPM [41], where ${\bm{P}}({\bm{x}}_{i}-{\bm{x}})=[1,({\bm{x}}_{i}-{\bm{x}})^{T}]^{T}$. Then (20) can reconstruct the function value $\hat{u}$ and provide an estimation of the gradient $\nabla\hat{u}$ at ${\bm{x}}$ as follows:
$\begin{bmatrix}\hat{u}\\ \nabla\hat{u}\\ \end{bmatrix}={\bm{M}}^{-1}({\bm{x}}){\bm{Q}}^{T}({\bm{x}}){\bm{D}}({\bm{x}}){\bm{u}},$ (21)
where ${\bm{u}}=[u_{1},\ldots,u_{N}]^{T}$ is the stacked sampled values, ${\bm{Q}}({\bm{x}})=[{\bm{P}}({\bm{x}}_{1}-{\bm{x}}),\ldots,{\bm{P}}({\bm{x}}_{N}-{\bm{x}})]^{T}$ is the stacked basis for every sample, ${\bm{D}}({\bm{x}})$ is the diagonal sample weighting matrix with ${\bm{D}}_{i,i}({\bm{x}})=d({\bm{x}}_{i}-{\bm{x}})$, and ${\bm{M}}({\bm{x}})={\bm{Q}}({\bm{x}})^{T}{\bm{D}}({\bm{x}}){\bm{Q}}({\bm{x}})$. We adopt the linear basis.

### v2

# Unstructured Moving Least Squares Material Point Methods: A Stable Kernel Approach With Continuous Gradient Reconstruction on General Unstructured Tessellations

@@ Abstract
The Material Point Method (MPM) is a hybrid Eulerian Lagrangian simulation technique for solid mechanics with significant deformation. Structured background grids are commonly employed in the standard MPM, but they may give rise to several accuracy problems in handling complex geometries. When using (2D) unstructured triangular or (3D) tetrahedral background elements, however, significant challenges arise (e.g., cell-crossing error). Substantial numerical errors develop due to the inherent $\mathcal{C}^{0}$ continuity property of the interpolation function, which causes discontinuous gradients across element boundaries. Prior efforts in constructing $\mathcal{C}^{1}$ continuous interpolation functions have either not been adapted for unstructured grids or have only been applied to 2D triangular meshes. In this study, an Unstructured Moving Least Squares MPM (UMLS-MPM) is introduced to accommodate 2D and 3D simplex tessellation. The central idea is to incorporate a diminishing function into the sample weights of the MLS kernel, ensuring an analytically continuous velocity gradient estimation. Numerical analyses confirm the method’s capability in mitigating cell crossing inaccuracies and realizing expected convergence.
@@ 2.1 Governing Equations
For clarity, in the remainder of this paper, the terms "grids" or "grid nodes" will exclusively refer to regular background grid nodes. In contrast, "mesh nodes" will denote nodes in general, unstructured meshes. For simplicity, we do not change the common abbreviations such as Particle-To-Grid (P2G) and Grid-To-Particle (G2P).
In the standard implementation of the MPM, physical quantities such as mass and velocity are retained at material points and then projected onto background grid nodes for further computation. (6) is discretized on these nodes by the Finite Element Method (FEM) and then solved using either implicit or explicit time integration schemes. This article focuses on the explicit symplectic Euler time integration method. While the extension to implicit methods is possible and straightforward, it would be orthogonal to the contribution of the article.
@@ 2.2 Explicit MPM Pipeline
The explicit MPM pipeline in each time step has four main stages: (1) the transfer of material point quantities to the background nodes, known as Particle-To-Grid (P2G), (2) the computation of the system’s evolution on these background nodes, (3) the back-transfer of the evolved quantities to the material points, known as Grid-To-Particle (G2P), and (4) the execution of necessary post-processings, such as elastoplasticity return mapping and material hardening. Algorithm 1 presents an overview of the MPM pipeline, and the main stages are elaborated below.
@@
1:Determine material point-node connectivity, calculate kernel functions $w_{p,i}$
@@
3:          Nodal mass: $m_{i}=\sum_{p}\rho_{p}V_{p}w_{p,i}$
4:          Nodal momentum: ${\bm{p}}_{i}=\sum_{p}{\bm{v}}_{p}\rho_{p}V_{p}w_{p,i}$
@@
6:          Internal force: ${\bm{f}}_{i}^{\text{int}}=-\sum_{p}V_{p}\bm{\sigma}_{p}\nabla w_{p,i}$
7:          Gravity: ${\bm{f}}_{i}^{\text{ext}}=\sum_{p}w_{p,i}m_{p}\mathbf{g}_{p}$
@@
14:          Update point velocities: ${\bm{v}}_{p}^{\Delta t}={\bm{v}}_{p}+\Delta t\sum_{i}w_{p,i}\ddot{{\bm{x}}}_{i}$
15:          Update point positions: ${\bm{x}}_{p}^{\Delta t}=\sum_{i}w_{p,i}\tilde{{\bm{x}}}_{i}$
16:Update deformation gradient: ${\bm{F}}_{p}^{\Delta t}=\left({\bm{I}}+\sum_{i}(\tilde{{\bm{x}}}_{i}-{\bm{x}}_{i})(\nabla w_{p,i})^{T}\right){\bm{F}}_{p}$
@@
19:Enforce plasticity, reset background deformation, advance to next timestep
@@
In the MPM, material points are the Lagrangian particles that track the location of the continuum along with physical attributes such as mass, position, and velocity. To evolve the dynamics on the background grid or mesh nodes, an interpolation function—also known as a transfer kernel or simply a kernel—needs to be determined to relate the information from particles to their nearby active nodes. Generally, for a particle located at ${\bm{x}}_{p}$ and all surrounding nodes at ${\bm{x}}_{1},\dots,{\bm{x}}_{N}$, the stacked kernel values associating the two sides are:
${\bm{w}}_{p}=[w_{p,1},\dots,w_{p,N}]^{T}={\bm{w}}({\bm{z}};{\bm{x}}_{p},{\bm{x}}_{1},\dots,{\bm{x}}_{N})\bigg{|}_{{\bm{z}}={\bm{x}}_{p}}.$ (7)
Specifically, we include ${\bm{x}}_{1},\dots,{\bm{x}}_{N}$ in this definition because constraints, such as the partition of unity, can only be determined while considering all active neighbors. Nonetheless, the neighbors are implicitly detected by ${\bm{x}}_{p}$; for conciseness, we omit this implicit condition in the remaining part of this paper. The kernel supports transferring information between the nodes and any location ${\bm{z}}$ near ${\bm{x}}_{p}$; in most MPM works, the kernel is evaluated at the lagged ${\bm{x}}_{p}$ before the completion of P2G, i.e., we set ${\bm{z}}={\bm{x}}_{p}$ and keep the kernel unchanged for both P2G and G2P.
Following this, the stacked gradients of the kernels are obtained with respect to the spatial variable ${\bm{z}}$, then evaluated at ${\bm{x}}_{p}$:
${\bm{G}}_{p}=[{\bm{g}}_{p,1},\dots,{\bm{g}}_{p,N}]^{T}=\left[\nabla_{{\bm{z}}}{\bm{w}}({\bm{z}};{\bm{x}}_{p})\right]\bigg{|}_{{\bm{z}}={\bm{x}}_{p}}.$ (8)
In the explicit MPM framework, the lumped mass at each background node is defined as $m_{i}=\sum_{p}\rho_{p}V_{p}w_{p,i}$, where $\rho_{p}$ represents the density and $V_{p}$ the volume of each nearby particle. This definition facilitates the calculation of the background node momentum, expressed as
@@
where $\ddot{{\bm{x}}}_{i}$ is the acceleration of node $i$, and ${\bm{f}}^{\text{int}}_{i}$ and ${\bm{f}}^{\text{ext}}_{i}$ represent the internal forces and the external forces acting on the it, respectively:
${\bm{f}}^{\text{int}}_{i}$ $=-\sum_{p}V_{p}\bm{\sigma}_{p}\nabla w_{p,i},$ (10)
${\bm{f}}^{\text{ext}}_{i}$ $=\sum_{p}m_{p}w_{p,i}{\bm{b}}_{p}+\sum_{p}m_{p}w_{p,i}{\bm{g}}_{p}.$ (11)
The stress tensor $\bm{\sigma}$ is determined by the deformation gradient ${\bm{F}}$ through some constitutive relation, indicating how material deformation influences internal forces:
$\bm{\sigma}=C({\bm{F}}).$ (12)
Stage 2: Evolution on the Background Nodes
Using the accelerations obtained from (9), we integrate the velocities and positions of the background nodes using a symplectic Euler time integrator employed throughout this work:
@@
where the time step size $\Delta t$ is chosen based on the CFL condition [18].
@@
The FLIP scheme [6] is utilized for all experiments discussed in Section 3. In FLIP, the particle positions and velocities are updated as follows:
$x_{p}^{\Delta t}$ $=\sum_{i}w_{p,i}\tilde{{\bm{x}}}_{i},$ (15)
$v_{p}^{\Delta t}$ $=v_{p}+\Delta t\sum_{i}w_{p,i}\ddot{{\bm{x}}}_{i}.$ (16)
Subsequently, the evolution of the deformation gradient ${\bm{F}}$ in (3) is conducted as follows:
${\bm{F}}_{p}^{\Delta t}=\left({\bm{I}}+\sum_{i}(\tilde{{\bm{x}}}_{i}-{\bm{x}}_{i})(\nabla w_{p,i})^{T}\right){\bm{F}}_{p}.$ (17)
Given the initial deformation gradient ${\bm{F}}^{0}={\bm{I}}$ and initial volume $V^{0}_{p}$, particle volumes are updated as:
@@
Stage 4: Post-Processing and Resetting the Background Nodes
This stage encompasses all post-processing tasks such as plasticity return mapping and material hardening [51]. In the updated Lagrangian MPM, the grid is reset to a non-deformed state at the end of each timestep. This is achieved by keeping the grid or mesh constant while zeroing all information such as velocity and acceleration.
@@ 2.3 Transfer Kernel
In the MPM, the transfer kernel is crucial for relaying particle information to adjacent background nodes. Techniques such as the B-spline MPM [52] and GIMP [2] use a specific compact support function to smoothly influence nearby grid nodes, whereas methods like Moving Least Squares MPM (MLS-MPM) [30] determine the kernel implicitly, based on the proximity of nodes. However, both strategies follow a similar workflow, which involves for every particle: (1) identifying the set of nearby nodes, and (2) calculating the transfer kernel and gradient for every particle-node pair.
This section first introduces the general MLS reconstruction process and the application of MLS-MPM with a comprehensive linear polynomial basis. It is followed by a discussion on a naive extension of MLS-MPM to unstructured meshes, highlighting the steps of identifying nearby nodes and computing the transfer weights. We then delve into the desirable properties of the kernel, emphasizing why the naive extension fails to yield continuous gradient reconstructions when particles cross cell boundaries. Finally, we propose a solution addressing the issue of discontinuous gradient reconstructions and introduce UMLS-MPM.
@@ 2.3.1 Introduction to General MLS and MLS-MPM
Given the kernel definition in (7), the Moving Least Squares (MLS) method aims to use a polynomial-based kernel to reconstruct $\hat{u}$ for some function $u$ at any location ${\bm{z}}$ near a given particle location ${\bm{x}}_{p}$. It is defined as follows:
$\hat{u}({\bm{z}};{\bm{x}}_{p})={\bm{p}}^{T}({\bm{z}}-{\bm{x}}_{p}){\bm{c}}({\bm{x}}_{p}),$ (19)
where ${\bm{p}}({\bm{z}}-{\bm{x}}_{p})=[p_{0}({\bm{z}}-{\bm{x}}_{p}),p_{1}({\bm{z}}-{\bm{x}}_{p}),\ldots,p_{l}({\bm{z}}-{\bm{x}}_{p})]^{T}$ represents the polynomial basis, ${\bm{c}}({\bm{x}}_{p})=[c_{0}({\bm{x}}_{p}),c_{1}({\bm{x}}_{p}),\ldots,c_{l}({\bm{x}}_{p})]^{T}$ are the corresponding coefficients, and $l$ indicates the total order of the basis. The coefficients ${\bm{c}}({\bm{x}}_{p})$ are determined by minimizing the sum of weighted square errors between the sampled function values $u_{i}$ and the reconstructed values $\hat{u}_{i}$ at nearby node positions ${\bm{x}}_{i}$:
${\bm{c}}({\bm{x}}_{p})=\mathop{\rm argmin}\sum_{i\in{\mathcal{B}}_{{\bm{x}}_{p}}}d({\bm{x}}_{i}-{\bm{x}}_{p})\,||u_{i}-{\bm{p}}^{T}({\bm{x}}_{i}-{\bm{x}}_{p}){\bm{c}}({\bm{x}}_{p})||^{2},$ (20)
where $d$ is a weighting function that takes proximity as input, and ${\mathcal{B}}_{{\bm{x}}_{p}}$ is the set of sample points in the local region around ${\bm{x}}_{p}$ where the weighting function is non-zero.
This minimization leads to the following solution for ${\bm{c}}({\bm{x}}_{p})$:
${\bm{c}}({\bm{x}}_{p})={\bm{M}}^{-1}({\bm{x}}_{p}){\bm{B}}({\bm{x}}_{p}){\bm{u}},$ (21)
where
${\bm{M}}({\bm{x}}_{p})$ $=\sum_{i\in{\mathcal{B}}_{{\bm{x}}_{p}}}d({\bm{x}}_{i}-{\bm{x}}_{p}){\bm{p}}({\bm{x}}_{i}-{\bm{x}}_{p}){\bm{p}}^{T}({\bm{x}}_{i}-{\bm{x}}_{p})$ (22)
$={\bm{P}}({\bm{x}}_{p}){\bm{D}}({\bm{x}}_{p}){\bm{P}}({\bm{x}}_{p})^{T},$
and
${\bm{B}}({\bm{x}}_{p})={\bm{P}}({\bm{x}}_{p}){\bm{D}}({\bm{x}}_{p}).$ (23)
Here we use the stacked notations: ${\bm{u}}=[u_{1},\ldots,u_{N}]^{T}$ is the stacked sample values, ${\bm{P}}({\bm{x}}_{p})=[{\bm{p}}({\bm{x}}_{1}-{\bm{x}}_{p}),\ldots,{\bm{p}}({\bm{x}}_{N}-{\bm{x}}_{p})]$ is the stacked basis, and ${\bm{D}}({\bm{x}}_{p})$ is the diagonal sample weighting matrix with ${\bm{D}}_{i,i}({\bm{x}}_{p})=d({\bm{x}}_{i}-{\bm{x}}_{p})$.
Substituting (21) into (19), we obtain the reconstruction:
$\hat{u}({\bm{z}};{\bm{x}}_{p})$ $={\bm{p}}^{T}({\bm{z}}-{\bm{x}}_{p}){\bm{M}}^{-1}({\bm{x}}_{p}){\bm{B}}({\bm{x}}_{p}){\bm{u}}$ (24)
$={\bm{w}}({\bm{z}};{\bm{x}}_{p})^{T}{\bm{u}},$
where the last derivation is obtained by defining the kernel for MLS, and note that ${\bm{M}}^{-1}({\bm{x}}_{p})$ is symmetric:
${\bm{w}}({\bm{z}};{\bm{x}}_{p})={\bm{B}}^{T}({\bm{x}}_{p}){\bm{M}}^{-1}({\bm{x}}_{p}){\bm{p}}({\bm{z}}-{\bm{x}}_{p}).$ (25)
Similar to (7) and (8), in the context of MPM, we again set the spatial variable ${\bm{z}}$ to be the particle positions ${\bm{x}}_{p}$ before completing P2G and obtain the kernel value:
${\bm{w}}_{p}={\bm{w}}({\bm{z}};{\bm{x}}_{p})\bigg{|}_{{\bm{z}}={\bm{x}}_{p}}={\bm{B}}^{T}({\bm{x}}_{p}){\bm{M}}^{-1}({\bm{x}}_{p}){\bm{p}}(0),$ (26)
as well as the kernel gradient:
${\bm{G}}_{p}=\left[\nabla_{{\bm{z}}}{\bm{w}}({\bm{z}};{\bm{x}}_{p})\right]\bigg{|}_{{\bm{z}}={\bm{x}}_{p}}={\bm{B}}^{T}({\bm{x}}_{p}){\bm{M}}^{-1}({\bm{x}}_{p})(\nabla_{{\bm{z}}}{\bm{p}})(0).$ (27)
@@
A special case involves using a complete linear polynomial basis, as in MLS-MPM [30], where ${\bm{p}}({\bm{z}}-{\bm{x}}_{p})=[1,({\bm{z}}-{\bm{x}}_{p})^{T}]^{T}$. Setting ${\bm{z}}={\bm{x}}_{p}$ in this basis, we have:
${\bm{p}}(0)$ $=\begin{bmatrix}1\\ \bm{0}_{\text{dim}}\end{bmatrix},$ (28)
$(\nabla_{{\bm{z}}}{\bm{p}})(0)$ $=\begin{bmatrix}\bm{0}^{T}_{\text{dim}}\\ {\bm{I}}_{\text{dim},\text{dim}}\end{bmatrix},$
$[{\bm{p}}(0),(\nabla_{{\bm{z}}}{\bm{p}})(0)]$ $={\bm{I}}_{\text{dim}+1,\text{dim}+1};$
Substituting (28) into (26) and (27) and stacking ${\bm{w}}_{p}$ and ${\bm{G}}_{p}$ in a column, we obtain a compact formula for both the kernel and the gradient:
$={\bm{B}}^{T}({\bm{x}}_{p}){\bm{M}}^{-1}({\bm{x}}_{p})[{\bm{p}}(0),(\nabla_{{\bm{z}}}{\bm{p}})(0)]$ (29)
$={\bm{B}}^{T}({\bm{x}}_{p}){\bm{M}}^{-1}({\bm{x}}_{p}){\bm{I}}_{\text{dim}+1,\text{dim}+1}$
$={\bm{B}}^{T}({\bm{x}}_{p}){\bm{M}}^{-1}({\bm{x}}_{p}).$
Applying (29) to the stacked function sample values at nodes ${\bm{u}}$, we obtain a compact formula for both the reconstructed function value $\hat{u}_{p}$ and the gradient $\nabla_{{\bm{z}}}\hat{u}_{p}$ of $u$:
$\begin{bmatrix}\hat{u}_{p}\\ \nabla_{{\bm{z}}}\hat{u}_{p}\\ \end{bmatrix}$ $=[{\bm{w}}_{p},{\bm{G}}_{p}]^{T}{\bm{u}}={\bm{M}}^{-1}({\bm{x}}_{p}){\bm{B}}({\bm{x}}_{p}){\bm{u}}.$ (30)
We adopt the linear basis throughout this work.